OpenStudy (anonymous):

What is the solution of -2x + 8 = 20

OpenStudy (lgbasallote):

-2x + 8 = 20 -2x = 20 - 8 -2x = 12 x = ?

OpenStudy (anonymous):

-6 thank you lgbasallote :))

OpenStudy (anonymous):

first you subtract 8 from both sides then you have -2x =12 then you divide by -2x from both sides and get x=-6

OpenStudy (lgbasallote):

\[\color{purple}{\textbf{<tips hat>}}\]